graph a line that goes through the point (4,4) and has a slope of -1/2
Answer by checkley71(8403)

Y=mX+b
4=-1/2*4+b
4=-2+b
b=4+2
b=6 THE Y INTERCEPT.
Y=-1/2X+6 IS THE LINE FORMULA